in young double slits experiment the ratio of intensities of a bright and dark bands are 6:1.the ratio of amplitude of interfering waves are
a. 16
b. 4
c. 5/3
d. ¼
 correct ans is option C but I want compelete solution....

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
I think you have put wrong details. I think ration should be 16 : 1 in question as only after that answer would be option C.
Imax / Imin = 16 / 1 = (a1 + a2)^2 / (a1 – a2)^2
a1 + a2 / a1 – a2 = 4/1
a1 / a2 = 5/3
 
Regards
Arun (askIITians forum expert)